OpenStudy (anonymous):

find volume of sphere w/ surface area of 144 sq in

OpenStudy (anonymous):

\[SA=4(\pi)r^2=144\] \[V=\frac{ 4 }{ 3 }\pi(r^3)\]

OpenStudy (anonymous):

surface area of a sphere is \(\large S=4 \pi r^2 \rightarrow r=\sqrt{\frac{S}{4\pi}} \) so volume is: \(\large V=\frac{4}{3}\pi r^3= \frac{4}{3}\pi (\sqrt{\frac{S}{4\pi}})^3\)

OpenStudy (anonymous):

thank you, lets see if i can work it.....